# Local MAP Sampling for Diffusion Models

### v1

@@ 1 Introduction
Optimization-based approaches—such as Resample (song2023solving, ), DiffPIR (zhu2023denoising, ), DCDP (li2024decoupled, ), and DMPlug (wang2024dmplug, )—have shown strong performance by alternating between denoising, optimization, and resampling to address inverse problems. Unlike DPS, which attempts to sample from the posterior distribution $p(x_{0}\mid y)$, optimization-based approaches prioritize reconstruction performance over distributional faithfulness. Nevertheless, their probabilistic foundation remains unclear: if the iterative procedure converges to the global MAP solution, i.e., $\arg\max p(x_{0}\mid y)$, would it still be consistent with DPS? Clarifying this foundation could provide both a principled interpretation and a stronger theoretical basis for optimization-based methods.
@@
  • Theoretical. We formulate Local MAP Sampling (LMAPS), a new inference framework that iteratively solves local maximum-a-posteriori subproblems along the diffusion trajectory. We analyze its relationship to global MAP and DPS, and show that LMAPS unifies Tweedie Moment Projected Diffusion (TMPD) and optimization-based inverse problem methods under a single framework. The relationship between LMAPS and existing methods are presented in Figure 1.
@@ 2 Background
Eq. 4 connects the conditional denoiser $\mathbb{E}[x_{0}\mid x_{t},y]$ with the unconditional denoiser $\mathbb{E}[x_{0}\mid x_{t}]$. However, the additional term $\nabla_{x_{t}}\log p(y\mid x_{t})$ is still intractable. One can train a neural network to approximate $\nabla_{x_{t}}\log p(y\mid x_{t})$, like classifier guidance (dhariwal2021diffusion, ). Training-free guidance, such as in (chung2022diffusion, ), usually approximates $\nabla_{x_{t}}\log p(y\mid x_{t})$ by a convenient single-sample approximation, $p(y\mid x_{t})\approx p(y\mid m_{0\mid t}(x_{t}))$, according to chain rule:
@@
Decoupled Annealing Posterior Sampling (DAPS) (zhang2025improving, ). Alternatively, DAPS developed a new framework to sample $x_{0}\sim\pi_{0\mid y}(x_{0}\mid y)$, which is given by the following iterations:
@@ 3.2 The difference between DPS, local MAP and global MAP
In summary, DPS targets $p(x_{0}\mid y)$, and LMAPS targets $\arg\max p(x_{0}\mid x_{t},y)$ at each step. Local MAP equals DPS only in Gaussian conditional settings; outside them, local MAP generally does not sample the posterior and can fail to reach the global MAP. We visualize a toy example in Figure 3. Compared to DPS and DAPS, LMAPS is less likely to generate samples in between-mode regions or low-density regions.
@@ 4 Local MAP sampling for inverse problem
The final ingredient for constructing a local posterior and solving the resulting MAP problem is the choice of prior $p(x_{0}\mid x_{t})$. While the true transition kernel of a diffusion model prior requires simulation, we can proceed as in previous work (boys2023tweedie, ; song2023pseudoinverse, ) by projecting onto the first two moments using a Gaussian approximation, $p(x_{0}\mid x_{t})\approx\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$, where $m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]$. While boys2023tweedie show that $\Sigma_{0\mid t}^{\textsc{tmpd}}(x_{t}):=\mathbb{E}[(x_{0}-m_{0\mid t})(x_{0}-m_{0\mid t})^{T}\mid x_{t}]=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0|t}$ to faithfully reflect the covariance of $x_{0}\mid x_{t}$, we will consider flexible choices of $\Sigma_{0|t}$. Finally, the local MAP problem amounts to solving
@@ 4.1 Approximated solution for nonlinear inverse problems
Relationship to optimization-based methods. Previous optimization-based approaches (song2023solving, ; li2024decoupled, ; zhu2023denoising, ) solve for $x_{0}^{*}$ through the following objective:
@@ 4.2 Exact Solution for Linear Inverse Problems
As discussed in Sec. 3.2, the local MAP solution matches the posterior mean for Gaussian posteriors $p(x_{t}\mid x_{0},y)$ arising from linear inverse problems $p(y\mid x_{0})=\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$ with a Gaussian assumption on the prior $p(x_{0}\mid x_{t})=\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$. Solving in closed form for the posterior mean as in (boys2023tweedie, ), we have
@@
We recover Tweedie Moment-Projected Diffusion (boys2023tweedie, ) as a special case for $\Sigma^{\textsc{tmpd}}_{0\mid t}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0\mid t}(x_{t})$, which is expensive since it requires the gradient with respect to the denoiser $m_{0\mid t}$.
@@ 5.1 Experimental setup
Inverse problems. We evaluate our method on image restoration and scientific inverse problems. For linear image restoration, we consider (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask). For nonlinear image restoration, we consider (1) phase retrieval, (2) high dynamic range (HDR) reconstruction, (3) nonlinear deblurring, (4) JPEG restoration, (5) quantization, where HDR, JPEG restoration and quantization are nonlinear inverse problems with non-differentiable operators. For scientific inverse problems, we adopt the benchmark from InverseBench (zheng2025inversebench, ), which includes Linear Inverse Scattering (LIS), Compressed sensing MRI (CS-MRI) and Black Hole Imaging. More details are provided in the App. C.
Dataset. For image restoration, we evaluated our method on FFHQ (karras2019style, ) $256\times 256$ and ImageNet $256\times 256$ datasets (deng2009imagenet, ). Following DAPS, we test the same subset of 100 images for both datasets. For scientific inverse problems, we adopt the same dataset as InverseBench (zheng2025inversebench, ).
Pretrained models. For image restoration tasks, we utilize the pre-trained checkpoint (chung2022diffusion, ) on the FFHQ dataset and the pre-trained checkpoint (dhariwal2021diffusion, ) on the ImageNet dataset. For scientific inverse problems, we adopt the pre-trained checkpoints from InverseBench.
@@
Metrics. For image restoration tasks, we report Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable, ). For scientific inverse problems, we primarily present PSNR in the main text, while additional task-specific metrics are provided in the App. D.
@@ 5.2 Main results
Ablation studies. Figure 4 presents ablation studies on optimization steps across different diffusion steps. The best performance is typically observed at NFE = 200–500, where increasing the number of optimization steps per diffusion step. yields notable improvements. Compared to the baseline SITCOM (600 NFEs with gradient computation through the U-Net), LMAPS attains similar performance while requiring substantially fewer computational resources. We report runtime comparisons for various methods on Deblurring task in Table 3 (App. B).
@@ 6 Related work
Recent advances in conditional generation have led to breakthroughs in text-to-image synthesis, semantic editing, and domain-specific applications such as image-to-image translation and controlled signal reconstruction (song2023loss, ; ye2024tfg, ; skreta2025feynman, ; singhal2025general, ; zheng2023guided, ). These methods have been especially impactful in solving inverse problems, including image restoration and scientific reconstruction tasks (wang2022zero, ; zheng2025inversebench, ). A wide range of approaches have been developed, spanning linear projection methods (wang2022zero, ; kawar2022denoising, ; zhang2025measurement, ; dou2024diffusion, ), Monte Carlo sampling (wu2023practical, ; phillips2024particle, ), variational inference (feng2023score, ; mardani2023variational, ; janati2024divide, ), and optimization-based strategies (song2023solving, ; zhu2023denoising, ; li2024decoupled, ; wang2024dmplug, ; alkhouri2024sitcom, ; he2023manifold, ).
Among these, Diffusion Posterior Sampling (DPS) and its variants (zhang2025improving, ; chung2022diffusion, ; song2023loss, ; yu2023freedom, ; rout2024beyond, ; yang2024guidance, ; bansal2023universal, ; boys2023tweedie, ; song2023pseudoinverse, ; ho2022classifier, ) have gained wide adoption due to their strong empirical performance and interpretability, as they directly sample from the posterior distribution $p(x_{0}\mid y)$. More recently, attention has shifted toward maximum a posteriori (MAP) estimation with diffusion priors. xu2025rethinking argued that DPS is in fact more consistent with the principle of MAP estimation rather than true posterior sampling, although their proposed sampling algorithm differs from ours. Finally, gutha2025inverse proposed sampling from the global MAP solution, $\arg\max p(x_{0}\mid y)$, though their approach is largely restricted to linear inverse problems.
@@ C.1 Dataset details
For scientific inverse problems, we adopt fluorescence microscopy images from InverseBench [40] on linear inverse scattering tasks, General Relativistic MagnetoHydroDynamic (GRMHD) [20] on black hole imaging, and multi-coil raw $k$-space data from the fastMRI knee dataset [36] on CS-MRI.
@@ C.2 Inverse problem details
Baselines from DAPS [37]. For image restoration tasks include: (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask), (6) phase retrieval, (7) high dynamic range (HDR) reconstruction, (8) nonlinear deblurring, we follow the same experimental setup as in DAPS.
InverseBench [40]. For scientific inverse problems, we adopt the setting introduced in InverseBench.
@@ C.3 Baseline details
For SITCOM [1], we use the hyperparameter configuration recommended in the original paper, with $N=20$ and $K=30$, resulting in $600$ NFEs and requiring gradient computation with respect to the U-Net.
@@
Other baselines we adopt the same reported results as in DAPS [37] and InverseBench [40].

### v2

@@ 1 Introduction
Optimization-based approaches—such as Resample (Song et al., 2023c), DiffPIR (Zhu et al., 2023), DCDP (Li et al., 2024), and DMPlug (Wang et al., 2024)—have shown strong performance by alternating between denoising, optimization, and resampling to address inverse problems. Unlike DPS, which attempts to sample from the posterior distribution $p(x_{0}\mid y)$, optimization-based approaches prioritize reconstruction performance over distributional faithfulness. Nevertheless, their probabilistic foundation remains unclear: if the iterative procedure converges to the global MAP solution, i.e., $\arg\max p(x_{0}\mid y)$, would it still be consistent with DPS? Clarifying this foundation could provide both a principled interpretation and a stronger theoretical basis for optimization-based methods.
@@
  • Theoretical. We formulate Local MAP Sampling (LMAPS), a new inference framework that iteratively solves local maximum-a-posteriori subproblems along the diffusion trajectory. We analyze its relationship to global MAP and DPS, and show that LMAPS unifies Tweedie Moment Projected Diffusion (TMPD) and optimization-based inverse problem methods under a single framework. The relationship between LMAPS and existing methods are presented in fig. 1.
@@ 2 Background
Eq. 4 connects the conditional denoiser $\mathbb{E}[x_{0}\mid x_{t},y]$ with the unconditional denoiser $\mathbb{E}[x_{0}\mid x_{t}]$. However, the additional term $\nabla_{x_{t}}\log p(y\mid x_{t})$ is still intractable. One can train a neural network to approximate $\nabla_{x_{t}}\log p(y\mid x_{t})$, like classifier guidance (Dhariwal and Nichol, 2021). Training-free guidance, such as in (Chung et al., 2022), usually approximates $\nabla_{x_{t}}\log p(y\mid x_{t})$ by a convenient single-sample approximation, $p(y\mid x_{t})\approx p(y\mid m_{0\mid t}(x_{t}))$, according to chain rule:
@@
Decoupled Annealing Posterior Sampling (DAPS) (Zhang et al., 2025a). Alternatively, DAPS developed a new framework to sample $x_{0}\sim\pi_{0\mid y}(x_{0}\mid y)$, which is given by the following iterations:
@@ 3.2 The difference between DPS, local MAP and global MAP
In summary, DPS targets $p(x_{0}\mid y)$, and LMAPS targets $\arg\max p(x_{0}\mid x_{t},y)$ at each step. Local MAP equals DPS only in Gaussian conditional settings; outside them, local MAP generally does not sample the posterior and can fail to reach the global MAP. We visualize a toy example in fig. 3. Compared to DPS and DAPS, LMAPS is less likely to generate samples in between-mode regions or low-density regions.
@@ 4 Local MAP sampling for inverse problem
The final ingredient for constructing a local posterior and solving the resulting MAP problem is the choice of prior $p(x_{0}\mid x_{t})$. While the true transition kernel of a diffusion model prior requires simulation, we can proceed as in previous work (Boys et al., 2023; Song et al., 2023b) by projecting onto the first two moments using a Gaussian approximation, $p(x_{0}\mid x_{t})\approx\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$, where $m_{0\mid t}(x_{t}):=\mathbb{E}[x_{0}\mid x_{t}]$. While Boys et al. (2023) show that $\Sigma_{0\mid t}^{\textsc{tmpd}}(x_{t}):=\mathbb{E}[(x_{0}-m_{0\mid t})(x_{0}-m_{0\mid t})^{T}\mid x_{t}]=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0|t}$ to faithfully reflect the covariance of $x_{0}\mid x_{t}$, we will consider flexible choices of $\Sigma_{0|t}$. Finally, the local MAP problem amounts to solving
@@ 4.1 Approximated solution for nonlinear inverse problems
Relationship to optimization-based methods. Previous optimization-based approaches (Song et al., 2023c; Li et al., 2024; Zhu et al., 2023) solve for $x_{0}^{*}$ through the following objective:
@@ 4.2 Exact Solution for Linear Inverse Problems
As discussed in Sec. 3.2, the local MAP solution matches the posterior mean for Gaussian posteriors $p(x_{t}\mid x_{0},y)$ arising from linear inverse problems $p(y\mid x_{0})=\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$ with a Gaussian assumption on the prior $p(x_{0}\mid x_{t})=\mathcal{N}(x_{0};m_{0\mid t},\Sigma_{0\mid t})$. Solving in closed form for the posterior mean as in (Boys et al., 2023), we have
@@
We recover Tweedie Moment-Projected Diffusion (Boys et al., 2023) as a special case for $\Sigma^{\textsc{tmpd}}_{0\mid t}=\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}m_{0\mid t}(x_{t})$, which is expensive since it requires the gradient with respect to the denoiser $m_{0\mid t}$.
@@ 5.1 Experimental setup
Inverse problems. We evaluate our method on image restoration and scientific inverse problems. For linear image restoration, we consider (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask). For nonlinear image restoration, we consider (1) phase retrieval, (2) high dynamic range (HDR) reconstruction, (3) nonlinear deblurring, (4) JPEG restoration, (5) quantization, where HDR, JPEG restoration and quantization are nonlinear inverse problems with non-differentiable operators. For scientific inverse problems, we adopt the benchmark from InverseBench (Zheng et al., 2025), which includes Linear Inverse Scattering (LIS), Compressed sensing MRI (CS-MRI) and Black Hole Imaging. More details are provided in the App. C.
Dataset. For image restoration, we evaluated our method on FFHQ (Karras et al., 2019) $256\times 256$ and ImageNet $256\times 256$ datasets (Deng et al., 2009). Following DAPS, we test the same subset of 100 images for both datasets. For scientific inverse problems, we adopt the same dataset as InverseBench (Zheng et al., 2025).
Pretrained models. For image restoration tasks, we utilize the pre-trained checkpoint (Chung et al., 2022) on the FFHQ dataset and the pre-trained checkpoint (Dhariwal and Nichol, 2021) on the ImageNet dataset. For scientific inverse problems, we adopt the pre-trained checkpoints from InverseBench.
@@
Metrics. For image restoration tasks, we report Peak Signal-to-Noise Ratio (PSNR), Structural SIMilarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018). For scientific inverse problems, we primarily present PSNR in the main text, while additional task-specific metrics are provided in the App. D.
@@ 5.2 Main results
Ablation studies. fig. 4 presents ablation studies on optimization steps across different diffusion steps. The best performance is typically observed at NFE = 200–500, where increasing the number of optimization steps per diffusion step. yields notable improvements. Compared to the baseline SITCOM (600 NFEs with gradient computation through the U-Net), LMAPS attains similar performance while requiring substantially fewer computational resources. We report runtime comparisons for various methods on Deblurring task in table 3 (App. B).
@@ 6 Related work
Recent advances in conditional generation have led to breakthroughs in text-to-image synthesis, semantic editing, and domain-specific applications such as image-to-image translation and controlled signal reconstruction (Song et al., 2023a; Ye et al., 2024; Skreta et al., 2025; Singhal et al., 2025; Zheng et al., 2023). These methods have been especially impactful in solving inverse problems, including image restoration and scientific reconstruction tasks (Wang et al., 2022; Zheng et al., 2025). A wide range of approaches have been developed, spanning linear projection methods (Wang et al., 2022; Kawar et al., 2022; Zhang et al., 2025b; Dou and Song, 2024), Monte Carlo sampling (Wu et al., 2023; Phillips et al., 2024), variational inference (Feng et al., 2023; Mardani et al., 2023; Janati et al., 2024), and optimization-based strategies (Song et al., 2023c; Zhu et al., 2023; Li et al., 2024; Wang et al., 2024; Alkhouri et al., 2024; He et al., 2023).
Among these, Diffusion Posterior Sampling (DPS) and its variants (Zhang et al., 2025a; Chung et al., 2022; Song et al., 2023a; Yu et al., 2023; Rout et al., 2024; Yang et al., 2024; Bansal et al., 2023; Boys et al., 2023; Song et al., 2023b; Ho and Salimans, 2022) have gained wide adoption due to their strong empirical performance and interpretability, as they directly sample from the posterior distribution $p(x_{0}\mid y)$. More recently, attention has shifted toward maximum a posteriori (MAP) estimation with diffusion priors. Xu et al. (2025) argued that DPS is in fact more consistent with the principle of MAP estimation rather than true posterior sampling, although their proposed sampling algorithm differs from ours. Finally, Gutha et al. (2025) proposed sampling from the global MAP solution, $\arg\max p(x_{0}\mid y)$, though their approach is largely restricted to linear inverse problems.
@@ C.1 Dataset details
For scientific inverse problems, we adopt fluorescence microscopy images from InverseBench [Zheng et al., 2025] on linear inverse scattering tasks, General Relativistic MagnetoHydroDynamic (GRMHD) [Mizuno, 2022] on black hole imaging, and multi-coil raw $k$-space data from the fastMRI knee dataset [Zbontar et al., 2018] on CS-MRI.
@@ C.2 Inverse problem details
Baselines from DAPS [Zhang et al., 2025a]. For image restoration tasks include: (1) super-resolution, (2) Gaussian deblurring, (3) motion deblurring, (4) inpainting (with a box mask), and (5) inpainting (with a 70% random mask), (6) phase retrieval, (7) high dynamic range (HDR) reconstruction, (8) nonlinear deblurring, we follow the same experimental setup as in DAPS.
InverseBench [Zheng et al., 2025]. For scientific inverse problems, we adopt the setting introduced in InverseBench.
@@ C.3 Baseline details
For SITCOM [Alkhouri et al., 2024], we use the hyperparameter configuration recommended in the original paper, with $N=20$ and $K=30$, resulting in $600$ NFEs and requiring gradient computation with respect to the U-Net.
@@
Other baselines we adopt the same reported results as in DAPS [Zhang et al., 2025a] and InverseBench [Zheng et al., 2025].